EXHIBIT 3: AMERICAN APPAREL INC. AND SUBSIDIARIES CONSOLIDATED BALANCE SHEET, 2008–2013

December 31,

(Amounts and shares in thousands, except per share amounts)

2013

2012

2011

2010

2009

2008

ASSETS

CURRENT ASSETS

Cash

$ 8,676

$ 12,853

$ 10,293

$ 7,656

$ 9,046

$ 11,368

Trade accounts receivable

20,701

22,962

20,939

16,688

16,907

16,439

Restricted cash

-

3,733

Prepaid expenses and other current assets

15,636

9,589

7,631

9,401

9,994

5,369

Inventories, net

169,378

174,229

185,764

178,052

141,235

148,154

Income taxes receivable and prepaid income taxes

306

530

5,955

4,114

4,494

604

Deferred income taxes, net of valuation allowance

599

494

148

626

4,627

3,935

Total current assets

215,296

224,390

230,730

216,537

186,303

185,869

Property and equipment, net

69,303

67,778

67,438

85,400

103,310

112,408

Deffered taxes

2,426

1,261

1,529

1,695

12,033

10,137

Other assets, net

46,727

34,783

25,024

24,318

25,933

25,195

TOTAL ASSETS

$ 333,752

$ 328,212

$ 324,721

$ 327,950

$ 327,579

$ 333,609

LIABILITIES AND STOCKHOLDERS' (DEFICIT) EQUITY

CURRENT LIABILITIES

Cash overdraft

$ 3,993

$-

$ 1,921

$ 3,328

$ 3,741

$ 2,413

Revolving credit facilities and current portion of long-term

debt

44,042

60,556

50,375

138,478

6,346

34,318

Accounts payable

38,290

38,160

33,920

31,534

19,705

32,731

Accrued expenses and other current liabilities

50,018

41,516

43,725

39,028

30,573

22,140

Fair value of warrant liability

20,954

17,241

9,633

993

2,608

8,582

Income taxes payable

1,742

2,137

2,445

230

Deferred income tax liability, current

1,241

296

150

-

Current portion of capital lease obligations

1,709

1,703

1,181

560

1,907

2,616

Total current liabilities

161,989

161,609

143,350

214,151

64,880

102,800

LONG-TERM DEBT, net of unamortized discount of $5,779 and

$27,929 at December 31, 2013 and 2012, respectively

213,468

110,012

97,142

444

Subordinated notes payble to related party

5,453

2,844

-

4,611

65,997

67,050

Capital lease obligations, net of current portion

1,726

542

4,355

3,292

Deferred tax liability

536

262

96

260

1,020

1,986

Deferred rent, net of current portion

18,225

20,706

22,231

24,924

22,052

16,011

Other long-term liabilities

11,485

10,695

12,046

7,994

11,934

6,058

Total Long -Term Liabilites

249,167

144,519

133,241

38,775

105,358

94,397

TOTAL LIABILITIES

411,156

306,128

276,591

252,926

$ 170,238

$ 197,197

STOCKHOLDERS' (DEFICIT) EQUITY

Shares outstanding at December 31, 2012

11

11

11

8

7

7

Additional paid-in capital

185,472

177,081

166,486

153,881

150,449

131,252

Accumulated other comprehensive loss

(4,306)

(2,725)

(3,356)

(3,168)

(2,083)

(2,703)

Accumulated deficit

(256,424)

(150,126)

(112,854)

(73,540)

19,012

17,900

Less: Treasury stock, 304 shares at cost

(2,157)

(2,157)

(2,157)

(2,157)

(10,044)

(10,044)

TOTAL STOCKHOLDERS' (DEFICIT) EQUITY

(77,404)

22,084

48,130

75,024

157,341

136,412

TOTAL LIABILITIES AND STOCKHOLDERS' (DEFICIT) EQUITY

$ 333,752

$ 328,212

$ 324,721

$ 327,950

$ 327,579

$ 333,609

Years Ended December 31 2012 2009 $ 633,941 $617,310S 547,336532,989 558,775 238,863 319,912 198,518 93,636 3,343 24,415 22,627 Net sales Cost of sales Gross profit Selling expenses General and administrative expenses Retail store impairment (Loss) income from operations Interest expense Foreign currency transaction loss Unrealized loss (gain) on change in fair value of warrants and purchase rights Loss (gain) on extinguishment of debt Other expense (income) Loss before income taxes Income tax provision Net loss Basic and diluted loss per share Weighted average basic and diluted shares outstanding Net loss (from above) Other comprehensive (oss) income item Foreign currency translation, net of tax Other comprehensive (oss) income, net of tax 313,056 320,885 241,683 106,957 289,927 327,383 227,447 97,327 1,647 962 41,559 120 252,436 294,900 209,841 104,085 4,267 (23,293) 33,167 1.679 253,080 279,909 218,198 103,167 8,597 (50,053) 23,752 (686) 1,540 (29,295) 39,286 (2,920) 4,126 (11,588) 204 (104.527) (33,459) (23,467) 993 (220) 4,928 39 (37,593) 1,771 12,164 1,721 (39,314) (86,315) $ (106,298) S (37,272) $ $ (0.96) $ (1.21) S (1.21) $ 71,626 0.01 71,026 1,112 105,980 $ (106,298) (37,272) 631 631 110,326 92,599 (39,314) (86,315) (188) 39 (1,085) (1,085) 87,400) $ 620 620 1,732 hensive loss $ (107,879) $ (36,641) $

EXHIBIT 4: CASH FLOW STATEMENTS FOR AMERICAN APPAREL, 2009-2013 in thousands) 2013 2012 2011 2010 2009 Net cash (used in) provided by Operating activities Investing activities Financing activities Effect of foreign exchange rate changes on cash Net (decrease) increase in cash $ (12,723) (25,147) 34,228 (535) S (4,177) $ 23,589 (24,853) 4,214 (390) $ 2,560 S 2,305 $(32,370) 45,203 (10,759) 15,662) 20,889) 48,172 (25,471) (1,165) S 2,637 1,390) (2,322) 12,582 (1,491) (1,530) Source: Compiled by the author from American Apparels 8-K reports for the year 2008-2013 American Apparel, Inc., Edgar Online, accessed March 27, 2016. EXHIBIT 5: AMERICAN APPARELS COMPARABLE STORE NUMBERS For the Quarter Ended Number 2013 0 stores 2011 Year March 31 June 30 September 30 December 31 235 238 241 238 243 249 237 244 248 237 242 244 2013

Use the information above calcuate the following financial ratios.

Current ratio:

Debt to Asset Ratio (Debt Ratio)

Inventory Turnover

Return on Sales (Profit Margin)
Answer #1

For the year 2013

Current ratio= Current assets/current liabilities

= 215296/161989

=1.33

Debt to Asset Ratio (Debt Ratio)

=Total debt/ Total assets

= 411156/333752

=1.23

Inventory Turnover= Cost of goods sold / Average inventory

= 313056/ ((169378+174229)/2)

=1.82

Return on Sales (Profit Margin)= Net Income/Sales

= -106298/633941

=-16.77%
Answer #2
1… Most Current Year(Dec. 2013) % Chg (+/–) Previous Year(Dec. 2012) 2… Analysis
Current Assets 215,296 -4.05% 224,390 Mainly due to decrease in inventory & cash
Total Assets 333752 1.69% 328212 Increase in other assets followed by Plant&equipment
Current Liabilities 161,989 0.24% 161,609 Quantum almost same in total,with compensating increases & decreases in its components.
Total Debt 249,167 72.41% 144,519 Mainly due to increase in long-term debt
Sales/Revenue 633941 2.69% 617310 Increase in revenue is a positive change that is good for the company
Cost of Goods Sold 313056 7.98% 289927 COGS has also increased as a % of sales from 47% in 2012 to 49% in 2013. Even though the increase is very small, it indicates rising cost to produce the items that are sold
Inventory 169,378 -2.78% 174,229 Decrease in inventory --may be due to more goods sold
Net Income/Loss -106298 185.20% -37272 Loss has increased indicating operational deficiency
3… Analysis
Current Ratio= Total Current assets/Total current liabilities 215296/161989= 1.33 Below the normal ratio level of 2--indicating that the company's liquidity position-ie. Capacity to meet its immediate obligations is not satisfactory.
Debt to Asset Ratio (Debt Ratio)=Total long-term liabilities/Total assets 249167/333752= 75% Indicates that assets are funded to the extent of 75% by long-term debt & balance by internally generated/own funds.
Inventory Turnover= COGS/Av.Inventory 313056/((174229+169378)/2)= 1.82 Inventory turns over at the rate of 1.82 times, ie. Once in (365/1.82= )200 days ,it is converted to sales .
Return on Sales (Profit Margin)=Net profit/Net sales -106298/633941= -16.77% The company is making a loss to the extent of 16.77% of sales revenue-- every $ 1 of sales, costs the company 1.1677